# A LLM-Powered Automatic Grading Framework with Human-Level Guidelines Optimization

### v1

@@ 5. Experiment
In this section, we conduct experiments to validate the effectiveness of GradeOpt. Through the experiments, we aim to answer the following research questions. (1) RQ1: Whether the refined guidelines based on prompt optimization match or exceed the performance of human-crafted guidelines? (2) RQ2: Are the optimized guidelines applicable to new datasets of the same or similar questions? (3) RQ3: How does each component contribute to the overall effectiveness of the guideline optimization system? Before presenting our experimental results and observations, we first introduce our general experimental settings.
5.1. Dataset
To answer the research questions above, we conduct experiments with a dataset of SAQs grading. Different from existing ASAG works (Dzikovska et al., 2013; Mohler et al., 2011), which focus on student work datasets, in this work we extend ASAG to a pedagogical answer grading task. These datasets collect answers from teachers over a set of questions designed to capture their content knowledge (CK) and pedagogical content knowledge (PCK) of mathematics (Copur-Gencturk et al., 2023). As the pedagogical answer grading requires a more intricate interpretation of the answer to understand the thought process, it will be valuable to explore the performance of GradeOpt in solving challenging ASAG tasks. Specifically, we selected three questions targeting teachers’ PCK and two questions targeting CK to test our framework. PCK helps test the working of our framework in capturing the thinking of responders, as PCK questions assess teachers’ ability to elaborate content to students and to reflect on students’ misconceptions. CK questions focus on examining teachers’ subject-related content knowledge through short answers, which is also a nontrivial task.
Overall, the pedagogical dataset is collected from two separate surveys finished by middle school mathematics teachers: one named pilot study denoted as $\mathcal{D}$, and the other named nationwide survey denoted as $\mathcal{D}^{\prime}$. The major difference between $\mathcal{D}$ and $\mathcal{D}^{\prime}$ is the scale and semantic complexity, as a large portion of $\mathcal{D}^{\prime}$ contains responses that only implicitly reflect the thought process, which makes them harder to distinguish. In our experiment, we select five questions that are shared by both $\mathcal{D}$ and $\mathcal{D}^{\prime}$. The data from $\mathcal{D}$ and $\mathcal{D}^{\prime}$ is used for the training stage and test-time adaptation stage, respectively. During the training stage, the answers to each question in $\mathcal{D}$ are split into training:validation:test with the ratio of 7:1:2. In the test-time adaptation stage, 100 samples from $\mathcal{D}^{\prime}$ are spared for test-time training and the remaining ones are used for test. The initial grading guidelines for each question are different, including types of assessment objectives and learning goals, which we refer to as expectations. Based on the number of expectations, we divide test questions into two types, i.e., single-expectation and multi-expectation. All answers are labelled using a three-point scale by two human raters. Detailed statistics of the datasets are included in Table 1.
Table 1. Detailed statistics of different questions in both Pilot Dataset $\mathcal{D}$ and National Dataset $\mathcal{D}^{\prime}$
Question ID Knowledge Type Pilot Dataset $|\mathcal{D}|$ National Dataset $|\mathcal{D^{\prime}}|$ Key Concept Multi-Expectation
$Q_{1}$ CK 261 1,352 $\times$ $\times$
$Q_{2}$ CK 265 1,364 $\times$ $\times$
$Q_{3}$ PCK 231 1,350
$Q_{4}$ PCK 233 1,348 $\times$
$Q_{5}$ PCK 228 1,127
5.2. Baselines
@@
We compared our model with several representative ASAG baselines. Firstly, we choose two popular non-LLM methods, i.e., SBERT (Reimers and Gurevych, 2019) with Logistic Regression and RoBERTa (Liu et al., 2019) with Fine-tuning. Both of them have demonstrated strong performance in prior studies (Condor et al., 2021; Poulton and Eliens, 2021). In addition, we adopt GPT-4o with zero-shot prompting as another baseline. Compared with non-LLM methods, LLM’s exceptional instruction and human-like reasoning capabilities make it a powerful method when facing complicated grading cases (Henkel et al., 2024). At last, to mitigate the manual burden of revising the guidelines in the GPT-4o zero-shot setting, we implement and compare GradeOpt with APO (Pryzant et al., 2023), which is a state-of-the-art method for automatic prompt optimization tasks.
5.3. Implementations
@@
To implement the nested iterative optimization, we set the outer batch size $|b_{out}|=64$ and inner batch size $|b_{in}|=8$. The outer loop iteration number $T=5$ and the inner loop iteration number $W=3$. We implement the beam search selection mechanism with Upper Confidence Bound (UCB) (Auer, 2003), where the guideline beam size $K=4$. The evaluation metric for UCB is Cohen’s Kappa as it empirically works better than other metrics. The agents in our framework are all powered by GPT-4o (OpenAI, 2024) with zero-shot prompting. The temperature for Grader is set to 0.0 to decrease the randomness of the result. The temperatures for both Reflector and Refiner are set to 0.5, since we want to encourage the LLMs to be more open in exploring the error reasons and propose the improving suggestions. For each question, we run the algorithm 3 times and report the average results.
5.4. Evaluation Metrics
In this work, we use Accuracy (Acc) and Cohen’s Kappa (Kappa) as the evaluation metrics to compare the performance of different models. To be specific, accuracy measures the percentage of correct predictions over all predictions and kappa measures the alignment between model’s and expert’s grading results.
5.5. Results on Pilot Dataset
In this section, we answer RQ1 with comparisons between baseline models and GradeOpt on the Pilot Dataset $\mathcal{D}$. From Table 2, we can observe that GPT-4o achieves the similar grading accuracy and Cohen’s Kappa comparing to traditional ML-based methods on the average result. This finding indicates that LLMs are a promising tool for ASAG task even without training on a domain dataset. Beyond that, it also reveals that LLMs have great potentials in providing generous solution to various ASAG tasks in the future. Additionally, using LLMs for ASAG provides the benefit of better comprehensibility as educators can easily see LLMs’ scoring rationale, which can be even given on the expectation level to assess the answer’s fulfillment of learning goals.
By comparing the results of APO and GradeOpt with GPT-4o, we can find both models bring steady performance gain to the original LLMs. This fact demonstrates our claim that using guidelines composed by pedagogical experts directly for LLMs is usually sub-optimal, which suggests that it is necessary to leverage additional guidance optimization algorithms to fully exploit the potential of LLMs in ASAG tasks. At last, our proposed GradeOpt always yields the best performance when compared to all baseline models over all the questions. This observation supports that GradeOpt can assist LLMs to achieve its optimal performance on ASAG tasks.
@@ 5.5. Results on Pilot Dataset
Table 2. Comparison of GradeOpt with Baseline Models on each question in Pilot Dataset $\mathcal{D}$. The best performed model of each metric is marked with bold, the second best one is marked with underline.
Model $Q_{1}$ $Q_{2}$ $Q_{3}$ $Q_{4}$ $Q_{5}$ Average
Acc Kappa Acc Kappa Acc Kappa Acc Kappa Acc Kappa Acc Kappa
RoBERTa 0.80 0.65 0.81 0.66 0.76 0.17 0.79 0.58 0.55 0.35 0.74 0.48
SBERT 0.61 0.32 0.70 0.42 0.76 0.17 0.69 0.44 0.70 0.42 0.69 0.35
GPT-4o 0.74 0.60 0.72 0.56 0.72 0.36 0.64 0.30 0.55 0.30 0.70 0.46
APO 0.90 0.85 0.89 0.80 0.76 0.39 0.79 0.60 0.76 0.63 0.82 0.65
GradeOPT 0.92 0.88 0.91 0.85 0.84 0.62 0.80 0.63 0.77 0.65 0.85 0.73
5.6. Results on National Dataset
We answer RQ2 by applying the optimized guidelines $G_{opt}$ learned from $\mathcal{D}$ in Section 5.5 to grade the test split of national dataset $\mathcal{D^{\prime}}_{test}$. To be specific, if the $G_{opt}$ fails to pass the OOD test, where the confidence indicator $\zeta<\mu$, we will implement the test-time training with the train split of national dataset $\mathcal{D^{\prime}}_{train}$. Then, the adapted guideline will be tested over the same $\mathcal{D^{\prime}}_{test}$ again. Overall, the results are reported in Table 3.
From the table, we can find that $Q_{1}$ is marked with OOD Flag as its confidence indicator $\zeta=-0.22<-0.2=\mu$. By comparing its performance between Table 2 and Tabel 3, we can confirm that it suffers great performance drops. Meanwhile, the confidence indicator of $Q_{2}$, $\zeta=-0.16>-0.2=\mu$, and its performance gap between $\mathcal{D}$ and $\mathcal{D}^{\prime}$ is relatively smaller. These two observations indicate that the proposed confidence indicator is a valid indicator for the OOD detection purpose. On the other hand, by comparing $G_{opt}$ with the raw guidelines provided by $G_{0}$, we find that $G_{opt}$ consistently outperforms $G_{0}$. This observation indicates that even the automatically optimized guideline suffers from the OOD issue, it is still better than the raw guideline. Finally, by calculating the performance change between the guideline before and after the test-time training, we can find that GradeOpt is able to get adapted to the new examples with only limited available annotated examples. In addition, the performance of guidelines after the test-time training is restored back to the acceptable grading range, e.g., Kappa$>0.6$, which indicates that the test-time training is an effective solution to help a high-performed $G_{opt}$ to quickly get applied to different datasets.
@@ 5.6. Results on National Dataset
Table 3. Test-time training result for Q1 and Q2 over National Dataset $\mathcal{D}^{\prime}$ before and after test-time training. The average confidence indicator (CI) on Pilot Dataset $\mathcal{D}$ is $\mu=-0.2$ and questions with CI ($\zeta<-0.2$) are marked with OOD.
Guideline Test-time Training $Q_{1}$ $Q_{2}$
Acc Kappa CI ($\zeta$) OOD Flag Acc Kappa CI ($\zeta$) OOD Flag
$G_{0}$ 0.67 0.49 - - 0.75 0.59 - -
$G_{opt}$$\times$ 0.70 0.52 -0.22 0.82 0.70 -0.16 $\times$
$G_{opt}$ 0.78 0.64 -0.17 $\times$ - - - -
Change - +0.08 +0.12 +0.05 - - - - -
5.7. Ablation Studies
To answer RQ3, we conduct ablation studies to the nested iteration introduced in Section 4.2. We choose to experiment with $Q_{4}$ as it is the only PCK question that contains one single expectation, as shown in Table 1. Our framework emphasizes PCK answers due to its challenge over CK answers in capturing the thinking process, as introduced in Section 1. The relatively simple rubric design of $Q_{4}$ makes the shortest guideline prompt, leaving room for GradeOpt to add in its reflective experience as it iteratively learns from $\mathcal{D}$. Thus, experimenting with $Q_{4}$ can better showcase GradeOpt’s optimization power. With the experimental results shown in the following sections, we demonstrate the effectiveness of each component.
First, we demonstrate the superiority of our misconfidence-based batch sampling strategy by comparing it with the random-based one. From Figure 6, we can observe that misconfidence-based batch sampling results are more consistent and accurate. While random selection generates optimal guidelines in 2 to 4 rounds, misconfidence-based selection consistently optimizes guidelines in 3 rounds. This, together with high predictive accuracy and alignment brought by misconfidence-based selection, makes the system reliable in practical educational scenarios as the required training round number is coherent.

### v2

@@ 5 Experiment
In this section, we conduct experiments to validate the effectiveness of GradeOpt. Through the experiments, we aim to answer the following research questions. RQ1: Whether the refined guidelines based on prompt optimization match or exceed the performance of human-crafted guidelines? RQ2: Are the optimized guidelines applicable to new datasets of the same or similar questions? RQ3: How does each component contribute to the overall effectiveness of the guideline optimization system?
5.1 Datasets
To address the research questions above, we conduct experiments using two representative datasets for SAQ grading. Unlike existing ASAG studies (dzikovska2013semeval, ; mohler2011learning, ), which focus solely on student responses, our work extends ASAG to the grading of pedagogical answers from both students and teachers. The first dataset, $\mathcal{D}_{1}$, consists of teachers’ responses to questions designed to assess the knowledge and skills essential for teaching mathematics (copur2022mathematics, ). Since grading pedagogical answers requires a more nuanced interpretation to capture the underlying thought process, evaluating GradeOpt on this dataset allows us to examine its performance on more complex ASAG tasks. Specifically, $\mathcal{D}_{1}$ includes six questions addressing different aspects of teacher knowledge, with responses labeled on a three-point scale: Bad (0), Fair (1), and Good (2). The second dataset, $\mathcal{D}_{2}$, evaluates GradeOpt on student responses, aligning with prior studies. It comprises 252 high school student responses to 11 assessment items within a physical sciences curriculum. These assessments measure Learning Progress (LP)-aligned scientific text-based explanations, reflecting students’ ability to apply knowledge of electrical interactions in high school Physical Science kaldaras2021developing . Responses in $\mathcal{D}_{2}$ are graded on a binary scale: Fail (0) or Pass (1). All grading labels in both $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ were assigned by at least two human raters. In cases of disagreement, a third rater provided the final judgment. Detailed statistics for both datasets are presented in Table 1. For our experiments, we split both datasets into training, validation, and test sets using a 7:1:2 ratio.
Table 1: Detailed statistics of different questions in both datasets $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. The number of samples in each label category is shown as $C_{i}$.
$\mathcal{D}_{1}$ $\mathcal{D}_{2}$
Question Total $C_{1}$ / $C_{2}$ / $C_{3}$ Question Total $C_{1}$ / $C_{2}$
$Q_{1}$ 261 36 / 104 / 121 $Q_{1}$ 252 43 / 209
$Q_{2}$ 265 78 / 47 / 140 $Q_{2}$ 252 123 / 129
$Q_{3}$ 236 132 / 66 / 38 $Q_{3}$ 252 113 / 139
$Q_{4}$ 231 180 / 44 / 7 $Q_{4}$ 252 183 / 69
$Q_{5}$ 232 83 / 112 / 37 $Q_{5}$ 252 242 / 10
$Q_{6}$ 229 74 / 43 / 112 $Q_{6}$ 252 244 / 8
$Q_{7}$ 230 64 / 114 / 52 $Q_{7}$ 252 245 / 7
$Q_{8}$ 231 108 / 24 / 99 $Q_{8}$ 252 227 / 25
- - - $Q_{9}$ 252 243 / 9
- - - $Q_{10}$ 252 210 / 42
- - - $Q_{11}$ 252 241 / 11
5.2 Baselines
@@
We compared our model with several representative ASAG baselines. Firstly, we choose two popular non-LLM methods, i.e., SBERT (reimers2019sentencebert, ) with Logistic Regression and RoBERTa (liu2019roberta, ) with Fine-tuning. Both of them have demonstrated strong performance in prior studies condor2021automaticSA ; poulton2021explaining . In addition, we adopt GPT-4o with zero-shot prompting, referred to as GPT-4o, as another baseline. Compared with non-LLM methods, LLM’s exceptional instruction and human-like reasoning capabilities make it a powerful method when facing complicated grading cases (henkel2024can, ). To mitigate the manual burden of revising the guidelines in the GPT-4o setting, we implement and compare GradeOpt with APO (pryzant2023gradient, ), which is a state-of-the-art method for automatic prompt optimization tasks.
5.3 Implementations
@@
To implement the nested iterative optimization, we set the outer batch size $|b_{out}|=64$ and inner batch size $|b_{in}|=8$. The outer loop iteration number $T=5$ and the inner loop iteration number $W=3$. We implement the beam search selection mechanism with Upper Confidence Bound (UCB) (auer2003using, ), where the guideline beam size $K=4$. The evaluation metric for UCB is Cohen’s Kappa as it empirically works better than other metrics. The agents in our framework are all powered by GPT-4o openai2024gpt4 with zero-shot prompting. The temperature for Grader is set to 0.0 to decrease the randomness of the result. The temperatures for both Reflector and Refiner are set to 0.5, since we want to encourage the LLMs to be more open in exploring the error reasons and propose the improving suggestions. For each question, we run the algorithm 3 times and report the average results.
5.4 Evaluation Metrics
In this work, we use Accuracy (Acc) and Cohen’s Kappa ($\kappa_{c}$) as the evaluation metrics to compare the performance of different models. To be specific, accuracy measures the percentage of correct predictions across all cases, while Cohen’s Kappa measures the inter-rater alignment between model’s predictions and expert annotations, accounting for agreement by chance. For the $\mathcal{D}_{1}$ dataset, which involves multi-class classification, we additionally utilize Quadratic Weighted Kappa ($\kappa_{w}$), which is particularly suitable for ordinal data as it assigns different weights to disagreements based on their magnitude.
5.5 Main Results
In this section, we address RQ1 by comparing baseline models with GradeOpt on both datasets, $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. Table 2 presents the performance of baseline models and GradeOpt on $\mathcal{D}_{1}$. The results reveal several key observations. While all models achieve relatively high accuracy across questions, the Cohen’s kappa values for baseline models such as RoBERTa and SBERT on some questions are notably low, often close to zero. This indicates a poor alignment between automated and manual grading. A deeper analysis reveals that non-LLM-based models exhibit a uniform majority classification phenomenon, leading to skewed grading patterns. This suggests that LLM-based models provide more reliable grading results compared to their non-LLM counterparts. Comparing GPT-4o with prompt-optimized methods further highlights the importance of optimization. Optimized prompts consistently enhance grading performance while reducing variance across different questions. This finding confirms that directly applying raw human-provided rubrics is suboptimal, and prompt optimization is necessary to fully leverage LLMs in automatic grading. Lastly, GradeOpt outperforms the state-of-the-art (SOTA) automatic prompt optimization method, APO, across all questions. This result demonstrates the superior effectiveness of GradeOpt in improving grading performance, reinforcing its advantage over existing methods.
@@ 5.5 Main Results
Table 2: Performance of GradeOpt and Baseline Models on $\mathcal{D}_{1}$. The best performed model of each metric is marked with bold, the second best one is marked with underline.
Question RoBERTa SBERT GPT-4o APO GradeOpt
Accuracy (Acc)
$Q_{1}$ 0.80 0.61 0.85 0.90 0.92
$Q_{2}$ 0.81 0.70 0.72 0.89 0.91
$Q_{3}$ 0.76 0.76 0.75 0.80 0.86
$Q_{4}$ 0.79 0.74 0.51 0.67 0.70
$Q_{5}$ 0.79 0.69 0.64 0.79 0.80
$Q_{6}$ 0.49 0.76 0.70 0.81 0.84
$Q_{7}$ 0.55 0.68 0.51 0.68 0.73
$Q_{8}$ 0.66 0.62 0.66 0.85 0.89
Cohen’s Kappa ($\kappa_{c}$)
$Q_{1}$ 0.65 0.32 0.76 0.85 0.88
$Q_{2}$ 0.66 0.42 0.56 0.80 0.85
$Q_{3}$ 0.00 0.00 0.38 0.51 0.68
$Q_{4}$ 0.00 0.00 0.09 0.35 0.36
$Q_{5}$ 0.58 0.44 0.30 0.60 0.63
$Q_{6}$ 0.00 0.17 0.55 0.69 0.70
$Q_{7}$ 0.00 0.41 0.33 0.50 0.52
$Q_{8}$ 0.37 0.29 0.48 0.75 0.80
Quadratic Weighted Kappa ($\kappa_{w}$)
$Q_{1}$ 0.70 0.31 0.82 0.88 0.89
$Q_{2}$ 0.81 0.52 0.75 0.93 0.94
$Q_{3}$ 0.00 0.00 0.61 0.67 0.76
$Q_{4}$ 0.00 0.00 0.24 0.41 0.54
$Q_{5}$ 0.59 0.32 0.56 0.68 0.71
$Q_{6}$ 0.00 0.17 0.62 0.77 0.80
$Q_{7}$ 0.00 0.48 0.58 0.64 0.70
$Q_{8}$ 0.44 0.33 0.68 0.84 0.87
Table 3: Performance of GradeOpt and Baseline Models on $\mathcal{D}_{2}$. The best performed model of each metric is marked with bold, the second best one is marked with underline.
Question RoBERTa SBERT GPT-4o APO GradeOPT
Accuracy (Acc)
$Q_{1}$ 0.84 0.84 0.84 0.96 0.98
$Q_{2}$ 0.57 0.55 0.76 0.86 0.86
$Q_{3}$ 0.69 0.67 0.88 0.92 0.92
$Q_{4}$ 0.65 0.65 0.78 0.80 0.82
$Q_{5}$ 0.94 0.94 0.94 0.94 0.98
$Q_{6}$ 0.98 0.98 0.88 0.92 0.94
$Q_{7}$ 0.98 0.98 1.00 1.00 1.00
$Q_{8}$ 0.92 0.92 0.47 0.76 0.78
$Q_{9}$ 0.98 0.98 0.96 0.96 0.98
$Q_{10}$ 0.88 0.88 0.73 0.86 0.88
$Q_{11}$ 0.94 0.94 0.90 0.96 0.96
Cohen’s Kappa ($\kappa_{c}$)
$Q_{1}$ 0.00 0.00 0.55 0.83 0.92
$Q_{2}$ 0.05 0.00 0.52 0.72 0.72
$Q_{3}$ 0.08 0.00 0.74 0.81 0.81
$Q_{4}$ 0.00 0.00 0.59 0.62 0.65
$Q_{5}$ 0.00 0.00 0.64 0.64 0.64
$Q_{6}$ 0.00 0.00 0.22 0.31 0.38
$Q_{7}$ 0.00 0.00 1.00 1.00 1.00
$Q_{8}$ 0.00 0.00 0.10 0.15 0.17
$Q_{9}$ 0.00 0.00 0.48 0.48 0.66
$Q_{10}$ 0.00 0.00 0.34 0.56 0.60
$Q_{11}$ 0.00 0.00 0.50 0.73 0.73
Table 3 presents the performance of baseline models and GradeOpt on dataset $\mathcal{D}_{2}$. Consistent with the findings in Table 2, LLM-based ASAG methods consistently outperform non-LLM-based models, reaffirming the advantages of leveraging LLMs for ASAG tasks. Moreover, GradeOpt consistently achieves the best performance among all baselines in Table 3, demonstrating its effectiveness across different datasets. This confirms that GradeOpt is a generalizable framework suitable for various ASAG tasks. Comparing model performance across the two datasets reveals an expected trend: grading responses related to teacher knowledge yields lower performance metrics. This aligns with our expectations, as evaluating pedagogical knowledge is inherently more complex, requiring deeper expertise and more intricate logical reasoning. Beyond performance improvements, LLM-based ASAG methods enhance grading transparency and comprehensibility. Educators can easily interpret the LLM’s scoring rationale, which can be further broken down at the expectation level. Each key concept or criterion specified in the rubric or learning objectives is individually assessed, allowing educators to evaluate how well student responses align with specific learning goals. This level of explainability fosters greater trust in automated grading and facilitates more informed instructional decisions.
5.6 Adaptation Results
@@
Table 4: Result for $Q_{1}$ and $Q_{2}$ over new dataset $\mathcal{D}_{1}^{\prime}$ before ($G_{opt}$) and after ($G_{opt}^{\prime}$) test-time training. The average confidence indicator (CI) on Pilot Dataset $\mathcal{D}_{1}$ is $\mu=-0.2$ and questions with CI ($\zeta<-0.2$) are marked with OOD.
Question Metric     $\mathbb{G_{0}}$     $\mathbb{G_{opt}}$     $\mathbb{G_{opt}^{\prime}}$     $\mathbb{\Delta}$
$Q_{1}$ Acc 0.67 0.7 0.78 +0.08
Kappa 0.49 0.52 0.64 +0.12
CI ($\zeta$) - -0.22 -0.17 +0.05
OOD - $\times$ -
$Q_{2}$ Acc 0.75 0.82 - -
Kappa 0.59 0.7 - -
CI ($\zeta$) - -0.16 - -
OOD - $\times$ - -
To address RQ2, we collect an external dataset, $\mathcal{D}_{1}^{\prime}$, containing responses to two questions, $Q_{1}$ and $Q_{2}$, from new teachers across the nation. Specifically, we gather 1,352 responses for $Q_{1}$ and 1,364 responses for $Q_{2}$. From these, we randomly select 100 responses per question to serve as the test-time training dataset, while the remaining responses are used for evaluation. By applying the optimized guidelines $G_{opt}$ learned from $\mathcal{D}_{1}$ in Section 5.5, we explore its grading performance over $\mathcal{D}_{1}^{\prime}$. In addition, if the $G_{opt}$ fails to pass the OOD test, where the confidence indicator $\zeta<\mu$, we will implement the test-time training with the train split of national dataset $\mathcal{D^{\prime}}_{train}$. Then, the adapted guideline will be tested over the same $\mathcal{D^{\prime}}_{test}$ again. Overall, the results are reported in Table 4. From the table, we can find that $Q_{1}$ is marked with OOD Flag as its confidence indicator $\zeta=-0.22<-0.2=\mu$. By comparing its performance between Table 2 and Tabel 4, we can confirm that it suffers great performance drops. Meanwhile, the confidence indicator of $Q_{2}$, $\zeta=-0.16>-0.2=\mu$, and its performance gap between $\mathcal{D}_{2}$ and $\mathcal{D}_{2}^{\prime}$ is relatively smaller. These two observations indicate that the proposed confidence indicator is a valid indicator for the OOD detection purpose. On the other hand, by comparing $G_{opt}$ with the raw guidelines provided by $G_{0}$, we find that $G_{opt}$ consistently outperforms $G_{0}$. This observation indicates that even the automatically optimized guideline suffers from the OOD issue, it is still better than the raw guideline. Finally, by calculating the performance change between the guideline before and after the test-time training, we can find that GradeOpt is able to get adapted to the new examples with only limited available annotated examples. In addition, the performance of guidelines after the test-time training is restored back to the acceptable grading range, e.g., Kappa $>0.6$, which indicates that the test-time training is an effective solution to help a high-performed $G_{opt}$ to quickly get applied to different datasets.
5.7 Ablation Studies
Side-by-side bars: misconfidence strategy accuracy vs. random.